# The SWE-Bench Illusion: When State-of-the-Art LLMs Remember Instead of Reason

### v1

@@ Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce a diagnostic task: file path identification from issue descriptions alone, to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions—without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
@@ 1 Introduction
To investigate this question systematically, we focus on a critical component of software engineering tasks: bug localization—the ability to identify which files contain bugs based on issue descriptions. This subtask is fundamental to successful issue resolution Hossain et al. (2024), yet should require genuine understanding of both the problem description and codebase structure. If models can accurately identify buggy file locations using only issue descriptions (without any repository context), this would suggest memorization of issue-file associations rather than reasoning ability.
@@
To systematically investigate these different memorization patterns, we design controlled experiments where models attempt to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances, despite lacking the contextual information that should be necessary for this task. Critically, performance drops by about 10% on tasks from the same repositories but not included in the SWE-Bench Verified, suggesting potential data contamination or benchmark-specific memorization rather than general coding proficiency.
To further investigate these patterns, we conduct cross-benchmark analysis comparing model performance on SWE-Bench against tasks from external repositories and different programming languages. Across ten evaluated LLMs, we consistently observe superior performance on SWE-Bench compared to external repositories, with particularly dramatic drops (up to 47 percentage points) when evaluating on repositories using different programming languages like C#. These findings raise serious concerns about the validity of current benchmark evaluations in software engineering and highlight the need for more robust, contamination-resistant evaluation approaches.
@@
    Diagnostic Task Design: We introduce a controlled bug localization task that isolates memorization effects from genuine problem-solving by asking LLMs to identify buggy files without providing repository context.
@@
    Empirical Evidence: We demonstrate through systematic experiments that current SWE-Bench performance may be inflated by memorization, with models achieving suspiciously high accuracy on context-free bug localization tasks.
@@
    Cross-Benchmark Analysis: We design a multi-benchmark validation process and provide evidence that performance disparities between SWE-Bench and external repositories suggesting repository-bias memorization rather than generalizable coding abilities, calling attention to evaluation validity concerns in the software engineering domain.
@@ 3.3 Metrics
We report accuracy as the primary metric to assess the model’s memorization of the task and its familiarity with the corresponding repository.
@@ 4.2 File Path Identification Accuracy on Filtered Instances
The persistence of these patterns in the filtered data is our most critical finding. It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns. This strongly suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
@@ 5.2 Benchmark Contamination
While existing mutation and metric-based methods are effective for single-function code generation, they primarily target instance-specific memorization and are less suited for complex, multi-file software engineering tasks like those in SWE-Bench. Our work introduces a complementary diagnostic approach that targets a different type of contamination: repository-bias memorization. Through subtask isolation—testing models on prerequisite tasks without necessary context—and systematic cross-benchmark analysis across different repository sets and programming languages, we provide a methodology to detect whether high benchmark performance reflects genuine problem-solving ability or disproportionate familiarity with specific codebases.
@@ 6 Conclusion
This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem, and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points) despite their popularity and likely training data inclusion. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
These findings suggest that widely-cited improvements in SWE-Bench performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities. The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills. Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than dataset-specific artifacts.

### v2

@@ Abstract
As large language models (LLMs) become increasingly capable and widely adopted, benchmarks play a central role in assessing their practical utility. For example, SWE-Bench Verified has emerged as a critical benchmark for evaluating LLMs’ software engineering abilities, particularly their aptitude for resolving real-world GitHub issues. Recent LLMs show impressive performance on SWE-Bench, leading to optimism about their capacity for complex coding tasks. However, current evaluation protocols may overstate these models’ true capabilities. It is crucial to distinguish LLMs’ generalizable problem-solving ability and other learned artifacts. In this work, we introduce two diagnostic task: file path identification from issue descriptions alone and ground truth function reproduction with only the current file context and issue description to probe models’ underlying knowledge. We present empirical evidence that performance gains on SWE-Bench-Verified may be partially driven by memorization rather than genuine problem-solving. We show that state-of-the-art models achieve up to 76% accuracy in identifying buggy file paths using only issue descriptions, without access to repository structure. This performance is merely up to 53% on tasks from repositories not included in SWE-Bench, pointing to possible data contamination or memorization. Similar patterns is also observed for function reproduction task, where the verbatim similarity is much higher on SWE-Bench Verified than on other similar coding benchmarks. These findings raise concerns about the validity of existing results and underscore the need for more robust, contamination-resistant benchmarks to reliably evaluate LLMs’ coding abilities.
@@ 1 Introduction
To investigate this question systematically, we focus on two critical components of software engineering tasks: bug localization, the ability to identify which files contain bugs based on issue descriptions, and patch generation, the ability to correct the buggy code. Both subtasks are fundamental to successful issue resolution. For bug localization Hossain et al. (2024), models should require genuine understanding of both the problem description and codebase structure to identify relevant files. For patch generation, models need to understand the semantics of the code and, potentially with additional context from the issue description, reason to generate the correct solution. Both capabilities could be compromised by memorization: models might recall specific issue-file associations or reproduce memorized code patterns rather than demonstrate genuine problem-solving ability.
However, directly measuring memorization versus genuine reasoning is challenging, as models legitimately learn coding patterns and domain knowledge during training. To address this challenge, we propose a differential testing approach inspired by software engineering practices: when ground truth oracles are unavailable, we use comparative analysis across similar systems as proxies. In our case, we evaluate models across multiple similar benchmarks and repositories, using performance disparities as indicators of memorization. Genuine coding ability should produce relatively consistent performance across comparable tasks, while memorization effects would manifest as systematic performance advantages on specific benchmarks where training data overlap is more likely.
@@
To systematically investigate these different memorization patterns, we design a controlled experiment pipeline that includes two diagnostic tasks: file-path identification and function reproduction. The file-path identification task requires models to identify buggy files using only GitHub issue descriptions, deliberately withholding all repository structure and code context. The function reproduction task measures verbatim similarity between generated patches and ground truth solutions. Our results reveal concerning patterns: state-of-the-art models like o3-mini OpenAI (2025b) achieve up to 76% accuracy on SWE-Bench-Verified instances for file-path identification, despite lacking the contextual information that should be necessary for this task. More critically, we observe substantial performance drops when evaluating on external benchmarks and repositories, with both file-path accuracy and verbatim similarity showing significantly higher values on SWE-Bench Verified compared to other evaluation sets, suggesting benchmark-specific memorization rather than generalizable coding proficiency.
These findings raise serious concerns about the validity of current benchmark evaluations in software engineering and highlight the need for more robust, contamination-resistant evaluation approaches.
@@
    Differential Testing for Memorization Detection: We introduce a systematic approach using performance disparities across similar benchmarks as memorization proxies when direct measurement is impossible.
@@
    Evidence of Benchmark Compromise: We demonstrate that state-of-the-art models achieve suspiciously high performance on SWE-Bench (76% context-free accuracy, elevated 5-gram similarity) with substantial drops on external benchmarks, indicating memorization rather than genuine capability.
@@
    Automated Evaluation Framework: We provide a reusable protocol for detecting memorization in any coding benchmark by comparing performance across similar evaluation sets.
@@ DISCUSSION
2.2 Function Reproduction Task
For the function reproduction task, we filter instances to include only function-level modifications and present models with issue descriptions alongside pre-patched files where modified functions are completely removed. Models must generate complete implementations for each missing function. Since function specification is unavailable, successful reproduction indicates memorization of specific textual sequences rather than contextual reasoning. We compare performance across multiple benchmarks to isolate memorization effects from general programming ability. We show the prompt template we used for function reproduction task in Appendix (Figure9).
Our approach directly tests memorization by requiring models to reproduce complete function implementations without access to signatures, docstrings, or relevant context from other files. This constraint ensures that successful reproduction requires prior exposure to the specific implementation rather than general problem-solving ability. The high consecutive 5-gram accuracy on SWE-Bench Verified compared to functionally equivalent control tasks provides evidence of memorized textual patterns rather than emergent programming competence.
@@ 3.3 Metrics
In this section, we details our choice of metrics used for our experiments.
@@
5-gram Consecutive Accuracy (5-gram Acc.)
To quantify the degree of verbatim memorization in model outputs, we compute 5-gram consecutive accuracy, which measures the proportion of predicted 5-grams (consecutive sequences of 5 tokens) that exactly match those found in the ground truth. This metric is calculated as the ratio of matching 5-grams to the total number of 5-grams in the prediction, with frequency-aware matching to prevent inflation from repeated phrases. Given that random 5-gram matches are statistically improbable, high scores on this metric provide strong evidence of memorization rather than genuine code generation. This allows us to distinguish models that have memorized data patterns in the training samples.
@@ 4.2 File Path Identification Accuracy on Filtered Instances
The persistence of these patterns in the filtered data is our most critical finding. It shows that the models’ ability to identify correct files is not due to simple text matching but is based on memorized knowledge, which includes instance-specific patterns and repository-biased patterns. This suggests that current benchmark scores are inflated by these memorization effects and do not purely reflect a model’s transferable software engineering skills.
Figure 6: Comparison of models’ consecutive 5-gram accuracy on function reproduction across benchmarks. Red ticks indicate improved verbatim similarity with the ground truth code, while blue ticks indicate a decrease.
4.3 Function Reproduction Result
We analyze model performance on the function reproduction task in this subsection. As shown in Figure 6, models achieve up to 35.0% consecutive 5-gram accuracy reproducing ground truth functions (the lower plot) from SWE-Bench Verified compared to up to 18.2% on instances outside the SWE-Bench ecosystem666 Due to limited resources, we only conducted experiments for o3-mini for the OpenAI reasoning models. . With limited context provided, only function names and issue descriptions, models lack sufficient information to systematically derive correct implementations. High consecutive 5-gram accuracy under these constraints requires reproducing exact textual sequences, indicating reliance on memorized training data rather than reasoning.
Instance-Specific Memorization
: Performance reveals a clear hierarchy that mirrors benchmark curation processes: SWE Verified (35.0%) >>> SWE Full (28.9%) >>> SWE Extra (18.3%). Critically, SWE Extra performance drops to the same level as external benchmarks (RefactorBench: 18.2%, Outside Tasks: 17.7%), despite originating from identical repositories as SWE Verified. This pattern indicates that models have specifically memorized the curated, canonical solutions that human evaluators deemed high-quality, rather than developing general familiarity with repository codebases.
These results demonstrate that SWE-Bench Verified performance substantially reflects memorization of training sequences rather than coding competence. The inability to reproduce functions from less-curated instances within the same repositories, combined with performance equivalence between SWE Extra and external benchmarks, reveals that apparent programming ability is largely an artifact of exposure to the specific, high-quality solutions selected for benchmark inclusion.
@@ 5.2 Benchmark Contamination
While existing mutation and metric-based methods are effective for single-function code generation, they face limitations when applied to complex, multi-file software engineering benchmarks like SWE-Bench. Task mutation approaches require generating semantically equivalent variations of real-world GitHub issues, a significantly more challenging task than mutating isolated coding problems. Metric-based approaches using perplexity or NLL cannot be applied to commercial models due to lack of access to hidden states, while n-gram similarity becomes noisy for complex patches where any functionally correct solution will share substantial textual overlap with ground truth.
Our work introduces a novel differential testing framework that addresses these limitations by using comparative performance analysis across related benchmarks as memorization proxies. Rather than requiring ground truth knowledge of training data exposure, we systematically compare model performance across benchmark variants and external evaluations to detect suspicious performance patterns. This differential testing methodology provides a systematic, reusable framework for detecting memorization in any coding benchmark without requiring access to training data or model internals.
@@ 6 Conclusion
This work presents evidence that current evaluations of LLM coding capabilities may be compromised by benchmark contamination and data memorization. Our systematic evaluation across ten models reveals two distinct memorization patterns: instance-specific memorization, evidenced by graduated performance decay within the SWE-Bench ecosystem (in both file-path identification task and function reproduction task), and repository-bias memorization, demonstrated by substantial performance drops on external repositories (up to 47 percentage points for file-path identification task) despite their popularity and likely training data inclusion. The consistency of these patterns across different model families and vendors indicates systematic exposure patterns rather than isolated issues.
These findings suggest that the improvements in SWE-Bench Verified performance may partially reflect benchmark-specific optimization rather than genuine advances in coding capabilities. The field urgently needs more robust evaluation frameworks that incorporate temporal controls to prevent training data contamination, cross-repository validation to test generalization beyond familiar codebases, and systematic cross-benchmark analysis to distinguish between memorization and transferable skills. Our work highlights the importance of developing contamination-resistant benchmarks and more sophisticated evaluation methodologies to ensure that reported progress reflects genuine advances in software engineering capabilities rather than dataset-specific artifacts.
@@ A.2 Additional Results
You are provided with a code repository and an issue description. Your task is to implement the complete function bodies for the marked RESPONSE comments. These functions are necessary to resolve the issue described in the problem statement. Repository: scikit-learn/scikit-learn <issue> Issue description here </issue> ... HINT: Fixes in _tags.py: modification in get_tags === File: sklearn/utils/_tags.py === <file> ... #TODO: RESPONSE 1: ... </file> ... You must provide complete implementations for each RESPONSE marker. Each function should: 1. Be properly indented to match its position in the file 2. Include the complete function body including the function signature (def line) 3. Address the issue described above RESPONSE INSTRUCTIONS: RESPONSE 1: Complete implementation of get_tags in sklearn/utils/_tags.py IMPORTANT: - Provide the COMPLETE function implementation, including the ’def’ line and all decorators if any - Maintain correct Python indentation - Each RESPONSE should contain the entire function from decorators (if any) through the complete function body - Number your responses to match the RESPONSE markers in the code - Wrap each response in triple backticks (‘‘‘) Please think through the issue first, then provide your implementations. DISCUSSION: Analyze the issue and plan your implementations here RESPONSE 1: ‘‘‘python # Your implementation here ‘‘‘ Function Reproduction Task Prompt Template
Figure 9: The prompt template used to test models’ ability to reproduce ground truth function without function specification.